hey
i need some help finding the area of a parallelogram with a base 2.7cm and a height of 2cm.
thanks

Answer by bucky(2097)

The equation that is used to find the area (A) of a parallelogram says that the area equals
the base (B) times the height (H) which is:
.
A = B*H
.
Now all you have to do is substitute the numbers that you are given in the problem. The
base (B) is 2.7 cm and the height (H) is 2.0 cm. Substituting those numbers result in:
.
A = 2.7 * 2.0 = 5.4 cm^2
.
Hope this helps you become more familiar with the properties of parallelograms.